# Structure-preserving contrastive learning for spatial time series

### v3

@@ 4.2 Hyperparameters
bs 8 [8, 16, 32, 64]  [^4] Maximum bs does not exceed train size.
@@ 5.1 Multivariate time series classification
Table 4 displays the classification performance on spatial and non-spatial UEA datasets. Next to the averaged accuracy, we also include the loss values on test sets to offer more information. More detailed results can be found in Tables B.1 and B.2 in B, where we present the classification accuracy with different representation learning losses for each dataset. Then we use Table 4 to more specifically compare the relative improvements induced by adding a topology or graph-geometry preserving regulariser. The improvement is measured by the percentage of accuracy difference from the corresponding baseline performance.
@@
Datasets Method Acc. AUPRC ${\mathcal{L}}_{\text{CLT}}$ ${\mathcal{L}}_{\text{SP}}$
With spatial features (14) TS2Vec 0.848 0.872 2.943
@@
Without spatial features (14) TS2Vec 0.523 0.555 8.417
@@
With spatial features (14) Topo-TS2Vec -4.403 0.800 16.54
@@
Without spatial features (14) Topo-TS2Vec -5.263 8.852 50.00
@@
Table 5: Structure preservation evaluation over datasets with and without spatial features in the UEA archive.
@@
kNN Trust. Cont. MRRE dRMSE kNN Trust. Cont. MRRE dRMSE
With spatial features (14) TS2Vec 0.563 0.868 0.875 0.117 0.346 0.419 0.784 0.765 0.189 0.150
Topo-TS2Vec 0.569 0.873 0.878 0.114 0.344 0.418 0.783 0.764 0.190 0.154
GGeo-TS2Vec 0.569 0.873 0.881 0.114 0.341 0.418 0.781 0.762 0.190 0.157
SoftCLT 0.562 0.866 0.875 0.117 0.348 0.420 0.788 0.765 0.187 0.171
Topo-SoftCLT 0.564 0.869 0.877 0.115 0.344 0.421 0.784 0.767 0.188 0.153
GGeo-SoftCLT 0.571 0.875 0.883 0.111 0.337 0.425 0.790 0.768 0.185 0.149
Without spatial features (14) TS2Vec 0.423 0.820 0.835 0.150 0.304 0.362 0.767 0.767 0.252 0.197
Topo-TS2Vec 0.424 0.820 0.831 0.151 0.308 0.356 0.763 0.767 0.254 0.191
GGeo-TS2Vec 0.420 0.820 0.832 0.151 0.310 0.365 0.769 0.771 0.253 0.189
SoftCLT 0.432 0.820 0.835 0.148 0.312 0.354 0.763 0.764 0.252 0.197
Topo-SoftCLT 0.426 0.818 0.834 0.148 0.312 0.361 0.768 0.768 0.254 0.205
GGeo-SoftCLT 0.430 0.822 0.835 0.147 0.315 0.355 0.761 0.762 0.257 0.203
@@
The assessment of similarity preservation is presented in Table 5 at both local and global scales. Consistent with the task-specific evaluation, Table 5 shows that structure-preserving regularisation preserves more complete information on similarity relations. The improvements are generally more significant on datasets with spatial features, which makes it more evident that our proposed preservation suits spatial time series data better. Although the comparisons in Tables 4$\sim$5 indicate more notable improvements by preserving graph geometry than preserving topology, we have to note that this does not demonstrate the universal superiority of one over the others. Different datasets have different characteristics that benefit from preserving global or local structure, and domain knowledge is necessary to determine which could be more effective. We will discuss this more in Section 6.
@@ 5.2 Macroscopic and microscopic traffic prediction
In Table 6, we present the performance evaluation for both macroscopic and microscopic traffic prediction. This table shows consistent improvements by pretraining encoders with our methods. Notably, single contrastive learning (i.e., TS2Vec and SoftCLT) does not necessarily improve downstream prediction, whereas it does when used together with preserving certain similarity structures. Given that our comparisons are conducted through controlling random conditions, this result effectively shows the necessity of preserving structure when learning traffic interaction representations.
Table 6: Macroscopic and microscopic traffic prediction performance evaluation.
@@
EVar
min. FDE
MR0.5
MR1
MR2
No pretraining 2.850 5.912 5.910 84.788 0.640 59.253 12.161 0.744
TS2Vec 2.878 5.982 5.982 84.415 0.636 59.453 11.899 0.558
Topo-TS2Vec 2.862 5.915 5.908 84.798 0.634 58.144 11.761 0.737
GGeo-TS2Vec 2.887 5.981 5.978 84.436 0.636 58.289 11.747 0.737
SoftCLT 2.856 5.938 5.932 84.675 0.641 59.501 11.940 0.785
Topo-SoftCLT 2.850 5.882 5.881 84.937 0.640 58.626 12.044 0.820
GGeo-SoftCLT 2.834 5.879 5.878 84.954 0.652 60.638 13.249 0.723
Best improvement (%) 0.568 0.558 0.547 0.196 0.929 1.872 3.399 25.000
@@
kNN Cont. Trust. MRRE dRMSE kNN Cont. Trust. MRRE dRMSE
@@ 5.3 Training efficiency
Incorporating structure-preserving regularisation increases computational complexity, and consequently, training time. The magnitude of this increase depends on the data and model that are applied on. With Table 8, we quantify the additional time required for structure preservation and evaluate its impact across diverse model architectures. In prior experiments, we used Convolutional Neural Network (CNN) encoders for the classification task on UEA datasets, Dynamic Graph Convolution Network (DGCN, [1]) encoder for macroscopic traffic prediction, and VectorNet [51] encoder for microscopic traffic prediction. To obtain a more comprehensive evaluation, we include two more Recurrent Neural Network (RNN) models for macroscopic traffic prediction: Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) encoders, paired with simple linear decoders.
@@
CNN 11.94 1.00$\times$ 1.46$\times$ 2.35$\times$ 1.00$\times$ 1.46$\times$ 2.36$\times$
@@
VectorNet 122.93 1.00$\times$ 1.45$\times$ 1.16$\times$ 1.23$\times$ 1.69$\times$ 1.42$\times$
MacroTraffic DGCN 74.78 1.00$\times$ 1.29$\times$ 1.09$\times$ 1.02$\times$ 1.40$\times$ 1.37$\times$
LSTM 18.04 1.00$\times$ 1.49$\times$ 1.12$\times$ 1.09$\times$ 1.57$\times$ 1.22$\times$
GRU 16.49 1.00$\times$ 1.54$\times$ 1.14$\times$ 1.10$\times$ 1.61$\times$ 1.24$\times$
a Detailed results are referred to Tables B.3 and B.4 in B.
@@
Figure 1: Fine-tuning progress of models trained from scratch and pretrained with different losses in macroscopic traffic prediction. Values of the final performance are referred to Table 6 for DGCN, to Tables C.1 and C.2 in C for LSTM and GRU.
@@ 6 Discussion
The theoretical foundation and experimental results presented in this paper not only demonstrate evident improvements in downstream task performance but also reveal a critical bridge between contrastive learning and similarity structure preservation. In this section, we discuss these findings to guide method selection and interpret the observed performance improvements.
@@ 7 Conclusion
This paper presents a method for structure-preserving contrastive learning for spatial time series, where we propose a dynamic mechanism to adaptively balance contrastive learning and structure preservation. This method is generally applicable to time series data with spatial or geographical features, which are particularly abundant in transportation systems. Extensive experiments demonstrate that our methods can improve the SOTA performance, including for multivariate time series classification in various contexts and for traffic prediction at both macroscopic and microscopic scales. Notably, preserving the original similarity structure in the latent space is shown to be well aligned with and beneficial to contrastive learning for spatio-temporal data. This is crucial for effective representational learning from large-scale data, as it impacts a neural network’s ability to model the underlying conditional distribution.
In general, adding structure-preserving regularisation has a limited impact on the efficiency of self-supervised representation learning, while facilitating the convergence of downstream tasks. It can be computationally intensive when the time sequence is long; however, evident performance improvement makes it an acceptable price to pay for utilising the information embedded in time series data. In addition, our experiments suggest that higher similarity structure preservation may be a good indicator of more informative representations, highlighting that the structural information of similarities in spatio-temporal data remains yet to be exploited. Given that many real-world practices involve spatial time series, this study and future research based on it can be applied not only to traffic interactions, but also to any that can benefit from preserving specific structures in similarity relations.

### v4

@@ 4.2 Hyperparameters
bs 8 [8, 16, 32]  [^3] Maximum bs does not exceed train size.
@@ 5.1 Multivariate time series classification
Table 3 displays the classification performance on spatial and non-spatial UEA datasets. Next to the averaged accuracy, we also include the loss values on test sets to offer more information. More detailed results can be found in Tables A.1 and A.2 in A, where we present the classification accuracy with different representation learning losses for each dataset. Then we use Table 4 to more specifically compare the relative improvements induced by adding a topology or graph-geometry preserving regulariser. The improvement is measured by the percentage of accuracy difference from the corresponding baseline performance.
@@
Datasets Method Acc. (↑) AUPRC (↑) ${\mathcal{L}}_{\text{CLT}}$ ${\mathcal{L}}_{\text{SP}}$
With spatial
features (14)
TS2Vec 0.848 0.872 2.943
@@
Without spatial
features (14)
TS2Vec 0.523 0.555 8.417
@@
With spatial
features (14)
Topo-TS2Vec -4.403 0.800 16.54
@@
Without spatial
features (14)
Topo-TS2Vec -5.263 8.852 50.00
@@
Table 5: Structure preservation evaluation over datasets with and without spatial features in the UEA archive. The standard deviations are computed across datasets.
@@
kNN (↑) Trust. (↑) Cont. (↑) MRRE (↓) dRMSE (↓) kNN (↑) Trust. (↑) Cont. (↑) MRRE (↓) dRMSE (↓)
With
spatial
features
TS2Vec 0.563 $\pm$ 0.149 0.875 $\pm$ 0.082 0.868 $\pm$ 0.073 0.117 $\pm$ 0.08 0.346 $\pm$ 0.099 0.419 $\pm$ 0.149 0.765 $\pm$ 0.110 0.784 $\pm$ 0.126 0.189 $\pm$ 0.124 0.150 $\pm$ 0.065
Topo-TS2Vec 0.569 $\pm$ 0.151 0.878 $\pm$ 0.082 0.873 $\pm$ 0.075 0.114 $\pm$ 0.081 0.344 $\pm$ 0.100 0.418 $\pm$ 0.150 0.764 $\pm$ 0.110 0.783 $\pm$ 0.125 0.190 $\pm$ 0.126 0.154 $\pm$ 0.069
GGeo-TS2Vec 0.569 $\pm$ 0.154 0.881 $\pm$ 0.08 0.873 $\pm$ 0.076 0.114 $\pm$ 0.079 0.341 $\pm$ 0.096 0.418 $\pm$ 0.153 0.762 $\pm$ 0.113 0.781 $\pm$ 0.129 0.190 $\pm$ 0.127 0.157 $\pm$ 0.075
SoftCLT 0.562 $\pm$ 0.156 0.875 $\pm$ 0.078 0.866 $\pm$ 0.074 0.117 $\pm$ 0.079 0.348 $\pm$ 0.096 0.420 $\pm$ 0.153 0.765 $\pm$ 0.112 0.788 $\pm$ 0.123 0.187 $\pm$ 0.125 0.171 $\pm$ 0.103
Topo-SoftCLT 0.564 $\pm$ 0.156 0.877 $\pm$ 0.077 0.869 $\pm$ 0.074 0.115 $\pm$ 0.077 0.344 $\pm$ 0.097 0.421 $\pm$ 0.157 0.767 $\pm$ 0.114 0.784 $\pm$ 0.126 0.188 $\pm$ 0.128 0.153 $\pm$ 0.068
GGeo-SoftCLT 0.571 $\pm$ 0.150 0.883 $\pm$ 0.073 0.875 $\pm$ 0.067 0.111 $\pm$ 0.076 0.337 $\pm$ 0.091 0.425 $\pm$ 0.149 0.768 $\pm$ 0.110 0.790 $\pm$ 0.121 0.185 $\pm$ 0.125 0.149 $\pm$ 0.065
Without
spatial
features
TS2Vec 0.423 $\pm$ 0.125 0.835 $\pm$ 0.095 0.820 $\pm$ 0.105 0.150 $\pm$ 0.105 0.304 $\pm$ 0.162 0.362 $\pm$ 0.175 0.767 $\pm$ 0.132 0.767 $\pm$ 0.136 0.252 $\pm$ 0.151 0.197 $\pm$ 0.136
Topo-TS2Vec 0.424 $\pm$ 0.125 0.831 $\pm$ 0.100 0.820 $\pm$ 0.106 0.151 $\pm$ 0.105 0.308 $\pm$ 0.164 0.356 $\pm$ 0.176 0.767 $\pm$ 0.134 0.763 $\pm$ 0.139 0.254 $\pm$ 0.154 0.191 $\pm$ 0.120
GGeo-TS2Vec 0.420 $\pm$ 0.126 0.832 $\pm$ 0.094 0.820 $\pm$ 0.108 0.151 $\pm$ 0.105 0.310 $\pm$ 0.145 0.365 $\pm$ 0.176 0.771 $\pm$ 0.137 0.769 $\pm$ 0.136 0.253 $\pm$ 0.157 0.189 $\pm$ 0.125
SoftCLT 0.432 $\pm$ 0.127 0.835 $\pm$ 0.099 0.820 $\pm$ 0.106 0.148 $\pm$ 0.105 0.312 $\pm$ 0.162 0.354 $\pm$ 0.177 0.764 $\pm$ 0.134 0.763 $\pm$ 0.136 0.252 $\pm$ 0.151 0.197 $\pm$ 0.131
Topo-SoftCLT 0.426 $\pm$ 0.119 0.834 $\pm$ 0.095 0.818 $\pm$ 0.104 0.148 $\pm$ 0.102 0.312 $\pm$ 0.163 0.361 $\pm$ 0.181 0.768 $\pm$ 0.131 0.768 $\pm$ 0.132 0.254 $\pm$ 0.152 0.205 $\pm$ 0.123
GGeo-SoftCLT 0.430 $\pm$ 0.122 0.835 $\pm$ 0.095 0.822 $\pm$ 0.101 0.147 $\pm$ 0.101 0.315 $\pm$ 0.151 0.355 $\pm$ 0.174 0.762 $\pm$ 0.134 0.761 $\pm$ 0.136 0.257 $\pm$ 0.153 0.203 $\pm$ 0.132
@@
[…]
@@ 5.2 Macroscopic and microscopic traffic prediction
In Table 6, we present the performance evaluation for both macroscopic and microscopic traffic prediction. This table shows consistent improvements by pretraining encoders with our methods. Notably, single contrastive learning (i.e., TS2Vec and SoftCLT) does not necessarily improve downstream prediction, whereas it does when used together with preserving certain similarity structures. Given that our comparisons are conducted through controlling random conditions, this result effectively shows the necessity of preserving structure when learning traffic interaction representations. In addition, we plot polar heatmaps in Figure 2 to visualise the encoded latent representations for the sensors in macroscopic traffic prediction. These sensors are deployed along a ring road, thus adjacent sensors are expected to have similar states and representation patterns. The figure intuitively shows better preserved spatial-temporal relations by contrastive learning and structure preservation.
Table 6: Macroscopic and microscopic traffic prediction performance evaluation. Metrics are reported as mean $\pm$ standard deviation over 10 stratified folds of the test set.
@@
EVar (↑)
min. FDE (↓)
MR0.5 (↓)
MR1 (↓)
MR2 (↓)
No pretraining 2.850 $\pm$ 0.044 5.911 $\pm$ 0.100 5.909 $\pm$ 0.100 84.784 $\pm$ 0.314 0.640 $\pm$ 0.013 59.253 $\pm$ 1.099 12.161 $\pm$ 0.929 0.744 $\pm$ 0.256
TS2Vec 2.878 $\pm$ 0.046 5.981 $\pm$ 0.111 5.981 $\pm$ 0.111 84.412 $\pm$ 0.337 0.636 $\pm$ 0.008 59.453 $\pm$ 0.866 11.899 $\pm$ 0.721 0.558 $\pm$ 0.17
Topo-TS2Vec 2.862 $\pm$ 0.043 5.914 $\pm$ 0.099 5.907 $\pm$ 0.099 84.793 $\pm$ 0.306 0.634 $\pm$ 0.008 58.144 $\pm$ 1.014 11.761 $\pm$ 0.809 0.737 $\pm$ 0.252
GGeo-TS2Vec 2.887 $\pm$ 0.045 5.980 $\pm$ 0.107 5.977 $\pm$ 0.107 84.433 $\pm$ 0.326 0.636 $\pm$ 0.010 58.289 $\pm$ 1.007 11.747 $\pm$ 0.797 0.737 $\pm$ 0.239
SoftCLT 2.856 $\pm$ 0.043 5.937 $\pm$ 0.105 5.931 $\pm$ 0.104 84.670 $\pm$ 0.348 0.641 $\pm$ 0.018 59.501 $\pm$ 1.742 11.940 $\pm$ 0.998 0.785 $\pm$ 0.278
Topo-SoftCLT 2.850 $\pm$ 0.045 5.881 $\pm$ 0.111 5.880 $\pm$ 0.111 84.934 $\pm$ 0.327 0.640 $\pm$ 0.012 58.626 $\pm$ 0.995 12.043 $\pm$ 0.735 0.820 $\pm$ 0.249
GGeo-SoftCLT 2.834 $\pm$ 0.046 5.878 $\pm$ 0.116 5.877 $\pm$ 0.116 84.952 $\pm$ 0.344 0.652 $\pm$ 0.013 60.638 $\pm$ 0.596 13.249 $\pm$ 0.681 0.723 $\pm$ 0.231
Best improvement 0.568 0.563 0.552 0.198 0.929 1.872 3.401 24.998
@@
Figure 2: Encoded representations after training with different losses on the test set of the Macroscopic traffic prediction task.
@@
kNN (↑) Trust. (↑) Cont. (↑) MRRE (↓) dRMSE (↓) kNN (↑) Trust. (↑) Cont. (↑) MRRE (↓) dRMSE (↓)
@@ 5.3 Training efficiency
Incorporating structure-preserving regularisation increases computational complexity, and consequently, training time. The magnitude of this increase depends on the data and model that are applied on. With Table 8, we quantify the additional time required for structure preservation and evaluate its impact across diverse model architectures. In prior experiments, we used Convolutional Neural Network (CNN) encoders for the classification task on UEA datasets, Dynamic Graph Convolution Network (DGCN, [1]) encoder for macroscopic traffic prediction, and Hierarchical Graph Neural Network (HGNN, based on VectorNet [59, 60]) encoder for microscopic traffic prediction. To obtain a more comprehensive evaluation, we include two more Recurrent Neural Network (RNN) models for macroscopic traffic prediction: Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) encoders, paired with simple linear decoders.
@@
[…]
@@
[…]
MacroTraffic
[…]
a Detailed results are referred to Tables A.3 and A.4 in A.
@@
Figure 3: Fine-tuning progress of models trained from scratch and pretrained with different losses in macroscopic traffic prediction. Values of the final performance are referred to Table 6 for DGCN, to Tables B.1 and B.2 in B for LSTM and GRU.
@@ 6 Discussion
The theoretical foundation and experimental results presented in this paper not only demonstrate evident improvements in downstream task performance but also reveal a critical bridge between contrastive learning and similarity structure preservation. In this section, we discuss these findings to guide method selection, interpret the observed performance improvements, and remind potential failure modes.
6.1 Method selection
@@
6.2 Failure modes
While the proposed method demonstrates consistent improvements across a range of tasks, several limitations and potential failure modes need to be acknowledged. These limitations are instructive in identifying the contexts where the method is most effective and where caution is warranted.
One source of difficulty arises from the sensitivity to data scale and complexity. The additional cost by structure-preserving regularisation is acceptable in many cases, but the overhead becomes prohibitive when time series are extremely long . For example, datasets with thousands of time steps result in a large similarity graph that inflates the computation memory and time of the GGeo-loss. Meanwhile, for models that are already well aligned with data, such as the DGCN baseline in macroscopic traffic prediction, pretraining with our method may provide only marginal benefits. This suggests that the utility of structure preservation is most pronounced when the data involve complex similarity relations that base models cannot easily capture.
Another important consideration concerns the benefits by dynamic weighting compared to fixed weights. The dynamic weighting mechanism is introduced to adaptively balance the loss magnitudes of contrastive learning and structure preservation. This works particularly well when the loss components have a large magnitude gap, which biases parameter updates. However, dynamic weighting may not outperform fixed weighting when the magnitudes of loss components are balanced in nature. In such circumstances, the adaptive mechanism can oscillate excessively, delaying convergence or amplifying noise in the latent space. Table 9 compares model performance on 5 datasets in the UEA archive. The datasets are selected because of relatively large magnitude gaps between contrastive learning loss and structure preservation loss. Dynamic and fixed weighting yield comparable classification accuracies, but the dynamic mechanism consistently improves structure preservation metrics and achieve lower loss.
Table 9: Comparison of dynamic and fixed weighting performance with the UEA archive. By fixed weighting we let $\sigma_{\text{CLT}}=1$ and $\sigma_{\text{SP}}=1$, so that the weights for both contrastive learning and structure preservation are 0.5.
Dataset Method Acc. (↑) kNN (↑) Trust. (↑) Cont. (↑) MRRE (↓) dRMSE (↓) ${\mathcal{L}}_{\text{CLT}}$ ${\mathcal{L}}_{\text{SP}}$
ERing GGeo-SoftCLT (dynamic weighting) 0.871 $\pm$ 0.011 0.776 $\pm$ 0.004 0.954 $\pm$ 0.002 0.940 $\pm$ 0.003 0.047 $\pm$ 0.001 0.209 $\pm$ 0.007 0.652 $\pm$ 0.023 106.972 $\pm$ 15.521
GGeo-SoftCLT (fixed weighting) 0.876 $\pm$ 0.013 0.767 $\pm$ 0.003 0.947 $\pm$ 0.002 0.933 $\pm$ 0.002 0.05 $\pm$ 0.001 0.226 $\pm$ 0.004 0.602 $\pm$ 0.015 149.115 $\pm$ 15.738
Libras GGeo-TS2Vec (dynamic weighting) 0.863 $\pm$ 0.010 0.757 $\pm$ 0.002 0.945 $\pm$ 0.001 0.915 $\pm$ 0.001 0.091 $\pm$ 0.001 0.281 $\pm$ 0.004 4.724 $\pm$ 0.038 387.867 $\pm$ 47.765
GGeo-TS2Vec (fixed weighting) 0.864 $\pm$ 0.016 0.754 $\pm$ 0.002 0.943 $\pm$ 0.001 0.915 $\pm$ 0.001 0.091 $\pm$ 0.001 0.275 $\pm$ 0.004 4.818 $\pm$ 0.021 519.156 $\pm$ 33.440
NATOPS GGeo-SoftCLT (dynamic weighting) 0.921 $\pm$ 0.009 0.640 $\pm$ 0.004 0.903 $\pm$ 0.004 0.859 $\pm$ 0.004 0.108 $\pm$ 0.002 0.296 $\pm$ 0.004 0.450 $\pm$ 0.007 712.646 $\pm$ 39.097
GGeo-SoftCLT (fixed weighting) 0.920 $\pm$ 0.011 0.639 $\pm$ 0.003 0.902 $\pm$ 0.003 0.855 $\pm$ 0.003 0.109 $\pm$ 0.002 0.295 $\pm$ 0.004 0.447 $\pm$ 0.006 811.255 $\pm$ 55.068
StandWalkJump GGeo-TS2Vec (dynamic weighting) 0.273 $\pm$ 0.049 0.389 $\pm$ 0.007 0.947 $\pm$ 0.003 0.905 $\pm$ 0.004 0.058 $\pm$ 0.002 0.226 $\pm$ 0.032 12.592 $\pm$ 2.747 3826.480 $\pm$ 2231.619
GGeo-TS2Vec (fixed weighting) 0.300 $\pm$ 0.057 0.386 $\pm$ 0.006 0.945 $\pm$ 0.003 0.904 $\pm$ 0.004 0.059 $\pm$ 0.002 0.239 $\pm$ 0.034 12.821 $\pm$ 2.847 4173.859 $\pm$ 2471.534
UWaveGestureLibrary GGeo-TS2Vec (dynamic weighting) 0.855 $\pm$ 0.011 0.759 $\pm$ 0.009 0.950 $\pm$ 0.005 0.956 $\pm$ 0.006 0.03 $\pm$ 0.002 0.308 $\pm$ 0.007 1.098 $\pm$ 0.170 501.284 $\pm$ 47.744
GGeo-TS2Vec (fixed weighting) 0.853 $\pm$ 0.018 0.755 $\pm$ 0.009 0.948 $\pm$ 0.006 0.954 $\pm$ 0.006 0.031 $\pm$ 0.002 0.306 $\pm$ 0.007 1.102 $\pm$ 0.174 604.046 $\pm$ 67.679
Finally, dependency on hyperparameter tuning presents a limitation. The approach proposed in this paper requires careful selection of kernel width for Laplacian approximation, temperature parameters in the contrastive loss, and the learning rate for dynamic weights. Although we provide a scheme for hyperparameter tuning, improper parameters can lead to ineffective training.
@@ 7 Conclusion
This paper presents a method for structure-preserving contrastive learning for spatial time series, where we propose a dynamic mechanism to adaptively balance contrastive learning and structure preservation. The method is generally applicable to time series data with spatial or geographical features, which are particularly abundant in transportation systems. Extensive experiments demonstrate that our methods improve the state of the arts, including for multivariate time series classification in various contexts and for traffic prediction at both macroscopic and microscopic scales.
An important advantage of this method is its ability to enforce the preservation of similarity structures in the latent space, thereby aligning representation learning with the original data manifold. At the same time, the efficiency costs of structure-preserving regularisation are moderate, making the approach practical for many applications. Nonetheless, several limitations deserve mention. The benefits of the method become expensive when time series are extremely long, where computational costs can be prohibitive. Dynamic weighting does not always outperform fixed weighting, but is particularly helpful when the training dynamics are dominated by one loss component. Moreover, the method depends on careful hyperparameter tuning, especially for the graph-geometry-preserving regulariser. These limitations do not undermine the contributions of the work but rather highlight opportunities for future research, including the design of lightweight approximations for long sequences, more stable dynamic weighting schemes, and principled strategies for automatic hyperparameter selection.
Notably, preserving the original similarity structure in the latent space is shown to be well aligned with and beneficial to contrastive learning for spatio-temporal data. This is crucial for informative representational learning from large-scale data, as it impacts a neural network’s ability to model the underlying conditional distribution. Our experiments suggest that higher similarity structure preservation is a good indicator of more informative representations, highlighting that the structural information of similarities in spatio-temporal data remains yet to be exploited. In the context of increasingly large neural network models that involve diverse data modalities, we hope this study sheds light on more effective training of large models in transportation systems.